2. A group of developers is opening a health club near a new housing development. The health club – which will have exercise and workout equipment, basketball courts, swimming pools, an indoor walking/running track, and tennis courts – is one of the amenities the developers are building to attract new homebuyers. The annual fixed cost for the building, equipment, utilities, staff, and so on is $600,000 and annual variable costs are $145 per member for thing like water, towels, laundry, soap, shampoo, and other member services. (a) The developers want the health club to break even at the end of the first year from the opening (12 months in total) by attracting 400 customers (assuming all the customers keep their memberships from the opening for the whole year, and no customer is added during the period). What is the minimal monthly fee to achieve their plan? (b) When they decide the membership fee as $250 per month, under the same assumption in (a), how many members will the club need at the opening in order to break even?
Answer #1

Annual fixed cost = $600000

Annual variable cost per member = $145

A) Total number of members = 400

Number of months for break even = 12

Let us assume monthly fee to be x.

For break even,

Annual fixed cost + Annual variable cost = Annual revenue

Hence, 600000 + 145*400 = 12*x*400

600000 + 58000 = 4800x

658000 = 4800x

x = 658000/4800

x = $137.08

Hence monthly fee is $137.08

B) Membership fee per month = $250

Let us assume members required to break even = y

Hence 600000 + 145*y = 12*250*y

600000 = 3000y - 145y

600000 = 2855y

y = 600000/2855 = 210.15 = 211

Hence number of members are 211.